OpenStudy (anonymous):

will fan and medal if question is answered! Solve the following system of equations algebraically. Show your work on your written work page. If the answer is not a whole number, give your answer accurate to 2 decimal places. Write the points in order from left to right. (The one that has the leftmost x-value goes first) \begin{cases} y = 3x^{2}-24x+52 \\ y = -3x+22 \end{cases}

OpenStudy (xapproachesinfinity):

first thing you need to do is to replace y in the 1 equation by \(\large \it\color{maroon}{-3x+22 }\)

OpenStudy (xapproachesinfinity):

can you do that?

OpenStudy (anonymous):

3x^2-24x+52 = -3x +22 ?

OpenStudy (anonymous):

still there?

OpenStudy (anonymous):

Yes that's correct : 3x^2 -24x +52 +3x -22=0 3x^2 -21x +30= 0 factor out a 3 3(x^2-7x+10) 3(x-2)(x-5) x= 2, 5

OpenStudy (xapproachesinfinity):

Sorry! something went wrong with OP and i didn't see the notice that you replied but Study is here! she will help you out

OpenStudy (anonymous):

@xapproachesinfinity I'm sorry too! >.< It was kinda bad of me. Sorry!! Yes for some reason I'm missing some notifications too. :o

OpenStudy (anonymous):

hehe np! thanks for your help :D

OpenStudy (xapproachesinfinity):

No worries @study100 , nothing to be sorry about Anytime sinista!